Parth (parthkohli):

\[\sum_{r=0}^n (-1)^r \binom{n}{r} \frac{1 + r \ln(10)}{(1 + \ln(10^n))^r}\]

OpenStudy (pawanyadav):

Can't see the question [Math Processing Error]

Parth (parthkohli):

\[\sum_{r=0}^n (-1)^r \binom{n}{r} \frac{1 + r \ln(10)}{(1 + \ln(10^n))^r}\]

OpenStudy (shadowlegendx):

PK u just reminded me I have some of these to do. ._. ;-;

Parth (parthkohli):

hey ur alive

OpenStudy (kainui):

just woke up

Parth (parthkohli):

here's a nice little hint to the previous problem: binomial theorem for any index

Parth (parthkohli):

which simply put is the maclaurin expansion of \((1+x)^n\)

Parth (parthkohli):

4:20 am!!!

OpenStudy (anonymous):

With \[(1+x)^n=\sum_{k=0}^n\binom nkx^k~~\implies~~x\frac{d}{dx}(1+x)^n=nx(1+x)^{n-1}=\sum_{k=0}^nk\binom nk x^k\]you can set \(x=-\dfrac{1}{1+n\ln10}\) so that you get \[\begin{align*}S&=\sum_{r=0}^n (-1)^r \binom{n}{r} \frac{1 + r \ln(10)}{(1 + \ln(10^n))^r}\\[1ex] &=\left(1-\frac{1}{1+n\ln10}\right)^n-\frac{n\ln10}{1+n\ln10}\left(1-\frac{1}{1+n\ln10}\right)^{n-1}\\[1ex] &=\left(1-\frac{1}{1+n\ln10}\right)^n\left[1-\frac{n\ln10}{1+n\ln10}\left(\frac{1+n\ln10}{n\ln10}\right)\right]\\[1ex] &=0 \end{align*}\]

Parth (parthkohli):

Well done!